OpenStudy (anonymous):

2q^2+2q/10q^2+4q, simplify by removing factors of 1

OpenStudy (anonymous):

\[\frac{2q^2+2q}{10q^2+4q}= \frac{2q(q+1)}{2q(5q+2)}\] \[=\frac{(q+1)}{(5q+2)}\]